OpenStudy (howard-wolowitz):

Find the first three terms of the sequence:

OpenStudy (howard-wolowitz):

\[a _{n}=2n+1\]

OpenStudy (howard-wolowitz):

\[a _{1}=\]

OpenStudy (howard-wolowitz):

\[a _{2}=\]

OpenStudy (howard-wolowitz):

\[a _{3}=\]

OpenStudy (howard-wolowitz):

@SolomonZelman

OpenStudy (howard-wolowitz):

@pitamar

OpenStudy (anonymous):

For \(a_1\) it is \(2 \cdot 1 + 1 = 2 + 1 = 3\). See how \(n\) was substituted by \(1\) here? So for \(a_2\) you have to substitute \(n\) with \(2\). can you try?

OpenStudy (howard-wolowitz):

i can but not sure how it will go.. eheh

OpenStudy (anonymous):

Try =)

OpenStudy (howard-wolowitz):

ok i got 1

OpenStudy (anonymous):

Show me your way

OpenStudy (howard-wolowitz):

but it should probbaly be 3

OpenStudy (howard-wolowitz):

considering 3,5,7

OpenStudy (anonymous):

Try not guessing. Can you show me your way?

OpenStudy (howard-wolowitz):

no im not guessing im saying those are the correct

OpenStudy (howard-wolowitz):

first one is 3 then the second is 5 and the third is 7

OpenStudy (anonymous):

Right =)

OpenStudy (anonymous):

So is it all clear?

OpenStudy (howard-wolowitz):

awesome

OpenStudy (anonymous):

Ok, great. So let me know if you have troubles

OpenStudy (howard-wolowitz):

i do here

OpenStudy (howard-wolowitz):

well i make a new question kk